My question is if the roots of the equation ax^2 + bx + c = 0 are m and n then the value of mn^2 + m^2n + m will be 
  1. c(a-b)÷a^2 
  2. 0
  3. -bc÷a^2
  4. None of this
My question is correct and I have rechecked it please provide me the correct answer

Grade:11

1 Answers

Ansh Mahapatra
5 years ago
I think you have to calculate mn^2+m^2n+m+n
=mn(m+n)+m+m
=(m+n)(mn+1)
=(-b/2a)(1+c/a)
[SUM OF ROOTS=-b/a]
[PRODUCT OF ROOTS=c/a]
=-bc/a^2